OpenStudy (anonymous):

John and Martha each had a steak dinner last night at Fred's diner. The steaks were $12.95 apiece, and they each had a glass of wine at $6 apiece. The bill included an additional 8% tax. How much was the total of John and Martha's bill?

OpenStudy (anonymous):

3/4 * 16 = ?

OpenStudy (anonymous):

{2(12.95+6) } *108/100 =

OpenStudy (anonymous):

Twenty six and twelve is thirty eight, less a dime is $37.90. Thirty eight times eight is sixteen less than 320, about 304, to get $3.04 tax. This gives a total of about $40.94 plus or minus a penny.

OpenStudy (anonymous):

{2(12.95+6) } *108/100 = $40.932

OpenStudy (anonymous):

I would bet that most states round that up to $40.94.

OpenStudy (anonymous):

right ^_^